a man travels 100 miles North from point A and than 200 miles west how far from point A is the man?
Answer by unlockmath(1122)

Hello,
We need to assume he travels directly west, then we can use the
C^2=A^2+B^2 formula. This is used to find the hypotenuse of a right triangle.
So plug in the numbers which will look like this:
c^2=100^2+200^2 Do the math will give us: (Note c^2 represents the diagonal)
c^2=10000+40000 Add like terms gives us:
c^2=50000 Now square root both sides gives us:
c=223.61 miles (Rounded off to the nearest hundredth)
RJ Toftness
www.math-unlock.com